OpenStudy (anonymous):

Yani buys a certain brand of cereal that costs $10 per box. Yani changes to a super-saving brand of the same size. The equation shows the price, y, as a function of the number of boxes, x, for the new brand. y = 7x Part A: How many more dollars is the price of a box of Yani's original brand of cereal than the price of a box of the super-saving cereal? Show your work. Part B: How much money does Yani save each month with the change in cereal brand if he buys 5 cereal boxes each month? Show your work.

OpenStudy (anonymous):

@jim_thompson5910 can you help with this also please?

jimthompson5910 (jim_thompson5910):

how much does the other brand cost?

OpenStudy (anonymous):

$7

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so it is $3?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

so that's part A

jimthompson5910 (jim_thompson5910):

Part B: How much money does Yani save each month with the change in cereal brand if he buys 5 cereal boxes each month? Show your work.

OpenStudy (anonymous):

3*5=15

OpenStudy (anonymous):

$15

OpenStudy (anonymous):

woops.5 each month so 15*5=75

jimthompson5910 (jim_thompson5910):

each box she saves $3 (assuming she goes for the cheaper brand) she buys 5 boxes total savings = 3*5 = 15 dollars

OpenStudy (anonymous):

so it is 15?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

np